# Straightening the Ruler: Field-Level Inference of the BAO Scale with LEFTfield

### v1

@@ Acknowledgments
We would like to thank Andrija Kostić for useful discussions and help with the MCMC analysis. We would also like to thank Julien Lesgourgues, Şafak Çelik, Minh Nguyen, Ivana Nikolac and Julia Stadler for many helpful discussions.
Appendix A Data Analysis
@@ A.1 Correlation length
In this section, we describe in detail how we conducted the analysis of field-level chains. We follow the same procedure for both mocks. For each bias expansion used to perform the inference, we analyze three $\Lambda$ values. For each $\Lambda$, we run three chains. For one chain, we leverage our knowledge of the ground truth for the initial conditions and start sampling from there. The other two chains begin from random initial conditions, with a different seed used for each to ensure they start from a unique set of initial conditions. All three chains also had slightly different initial values for the cosmological parameters sampled. This approach ensured that the inferred value of BAO is independent of the sampler’s starting point. As mentioned in the Sec. 5.1, one chain used Gaussian prior for bias, while the other two used uniform priors; however, these priors were essentially uninformative, as the bias parameters are well constrained by the data (at least when sampling only $\beta$ as cosmological parameter),
To perform the analysis, we combine the three independent chains into one large chain, which consists of $N$ elements; for a parameter $f$, we label the sample points as $\{f_{s}\}$, after discarding the burn-in portion of each chain, which was about 5 correlation lengths. To denote the mean of the dataset, we use $\langle f_{s}\rangle\equiv\bar{f}=\frac{1}{N}\sum_{s=1}^{N}f_{s}$. We continue running the chains until we achieve 100 effective samples of the $\beta$ parameter for each dataset and $\Lambda$ considered. The number of effective samples is calculated by dividing the total number of samples, $N$, by the auto-correlation length. The auto-correlation length for an MCMC chain indicates the number of steps required within the chain to obtain samples that are independent of each other. Correlation length can be estimated using the relation
@@
is the auto-correlation function. $M$ is the maximum separation between the samples considered. As the best choice for $M$, [47] suggests using the smallest value of $M$ such that $M\geqslant C\hat{\tau}_{f}(M)$ for a constant $C$ chosen to minimize the covariance of the estimator. This is usually achieved for values of $C$ close to 5. Plots indicating $\tau_{f}$ values of $\beta$ for chains analyzed in the paper are shown below. In the tables below, we summarize the correlation lengths for the parameter $\beta$.
@@
Table 2: Correlation length values for chains run on the Mock A using Lagrangian bias.
@@
Table 3: Correlation length values for chains run on the Mock B using Lagrangian bias.
@@
Table 4: Correlation length values for chains run on the Mock B using Eulerian bias.
@@ A.2 Parameter posteriors: marginalized likelihood
Below we present the corner plots showing posterior distribution of parameters sampled for both mocks and cutoffs. Results for the Mock B are summarized in the Fig. 10, where again we denote Lagrangian bias results in blue and Eulerian in orange. We notice the same trend of increasing noise as we did in Sec. 5.2. In Fig. 11, we present the results for Mock A. The effective noise also grows in the case of Mock A, but in this case it is $\sigma_{\epsilon}$ that grows more significantly, rather than $\sigma_{\epsilon,2}$.
@@
Figure 10: Posterior of parameters for the inference performed on Mock B. Top two panels (orange color) corresponds to Eulerian bias, while the lower two panels (blue) correspond to Lagrangian bias. The intermediate cutoff $\Lambda=0.18\,h\,{\rm Mpc}^{-1}$ behaves similarly, and is not shown here for brevity.
@@
Figure 11: Posterior of parameters for the inference performed on Mock A.
@@ A.3 Parameter posteriors: non-marginalized likelihood
Finally, Figure 14 displays the corner plots for all sampled parameters, including the mean value and the size of the error bars for each parameter. Comparing the size of the error bars for $\sigma(\beta)$ in this plot with those in Figure 10(a) reveals that the error bar size for $\beta$ does not depend on the type of likelihood used, as expected. Furthermore, the values of $\sigma_{\epsilon}$ and $\sigma_{\epsilon,2}$ are consistent with those obtained using the marginalized likelihood.
@@
Figure 14: Results of the joint FreeIC inference on the Mock B using non-marginalized likelihood and Eulerian bias. We show the posterior for all noise parameters, bias parameters and $\beta$. Inference was performed at $\Lambda=0.15\,h\,{\rm Mpc}^{-1}$.
The same analysis using the non-marginalized likelihood has been applied to Mock A sampled with the Lagrangian bias. The results of this analysis are summarized in Fig. 15. Similarly as in the case of mock B, we find that the error bar $\sigma(\beta)$ is the same as when the marginalized likelihood is used for the same cutoff.
Figure 15: Results of the joint FreeIC inference on the Mock A using non-marginalized likelihood and Lagrangian bias. We show the posterior for all noise parameters, bias parameters and $\beta$. Inference was performed at $\Lambda=0.15\,h\,{\rm Mpc}^{-1}$.

### v2

@@ Acknowledgments
We would like to thank Andrija Kostić for his help with setting up the analysis of the MCMC chains, and Uroš Seljak for insightful discussions that led us to conduct more validation tests. We would also like to thank Julien Lesgourgues, Ariel Sánchez, Şafak Çelik, Minh Nguyen, Ivana Nikolac and Julia Stadler for many helpful discussions.
Appendix A Posterior diagnostics and detailed results
@@ A.1 Correlation length
In this section, we describe in detail how we conducted the analysis of field-level chains. We follow the same procedure for both mocks. For each data set, bias model, and cutoff, we ran three chains: one chain starts from the ground-truth initial conditions, while the other two chains begin from random initial conditions, with a different seed used for each to ensure they start from a distinct set of initial conditions. All three chains also have slightly different initial values for the model parameters that are being sampled. This approach ensures that the inferred value of BAO is independent of the sampler’s starting point. As mentioned in Sec. 5.1, one chain uses Gaussian priors for the bias parameters, while the other two uses uniform priors; however, both of these priors are essentially uninformative, as the bias parameters are well constrained by the data, as shown for the non-marginalized cases in Figs. 14–fig: Mock A Fourier.
To perform the analysis, we combine the three independent chains into one large chain, which consists of $N$ samples; for a parameter $f$, we label the sample points as $\{f_{s}\}$, after discarding the burn-in portion of each chain, which is about 5 correlation lengths. To denote the mean of the dataset, we use $\langle f_{s}\rangle\equiv\bar{f}=\frac{1}{N}\sum_{s=1}^{N}f_{s}$. We continue running the chains until we achieve at least 100 effective samples of the $\beta$ parameter for each dataset and $\Lambda$ considered. The number of effective samples is calculated by dividing the total number of samples, $N$, by the auto-correlation length. The auto-correlation length for an MCMC chain indicates the number of steps required within the chain to obtain samples that are independent of each other. Correlation length can be estimated using the relation
@@
is the auto-correlation function. $M$ is the maximum separation between the samples considered. As the best choice for $M$, Ref. [49] suggests using the smallest value of $M$ such that $M\geqslant C\hat{\tau}_{f}(M)$ for a constant $C$ chosen to minimize the covariance of the estimator. This is usually achieved for values of $C$ close to 5. Plots indicating $\tau_{f}$ values of $\beta$ for chains analyzed in the paper are shown in Fig. 2 and Fig. 13. In the tables below, we summarize the correlation lengths for the parameter $\beta$. There is significant scatter among the estimates for $\tau$, but all are significantly below 1,000.
@@
Table 4: Correlation length values for chains run on Mock A using Lagrangian bias.
@@
Table 5: Correlation length values for chains run on Mock B using Lagrangian bias.
@@
Table 6: Correlation length values for chains run on Mock B using Eulerian bias.
@@ A.2 Parameter posteriors: marginalized likelihood
Below we present the corner plots showing posterior distribution of parameters sampled for both mocks and cutoffs. Results for Mock B are summarized in Fig. 10, where again we denote Lagrangian bias results in blue and Eulerian in orange. We notice the same trend of increasing noise as we did in Sec. 5. In Fig. 11, we present the results for Mock A. The effective noise also grows in the case of Mock A, but in this case it is $\sigma_{\epsilon}$ that grows more significantly, rather than $\sigma_{\epsilon,2}$.
@@
Figure 10: Parameter posteriors for the inference performed on Mock B. Top two panels (orange color) correspond to Eulerian bias, while the lower two panels (blue) correspond to Lagrangian bias. The left panels show results for $\Lambda=0.15\,h\,{\rm Mpc}^{-1}$, while the right panels those for $\Lambda=0.2\,h\,{\rm Mpc}^{-1}$. The intermediate cutoff $\Lambda=0.18\,h\,{\rm Mpc}^{-1}$ behaves similarly, and is not shown here for brevity.
@@
Figure 11: Parameter posteriors for the inference performed on Mock A.
@@ A.3 Parameter posteriors: non-marginalized likelihood
Finally, Figure 14 displays the corner plots for all sampled parameters, including the mean value and the size of the error bars for each parameter. Comparing the size of the error bars for $\sigma(\beta)$ in this plot with those in Figure 10(a) reveals that the error bar size for $\beta$ does not depend on whether bias parameters are sampled or marginalized analytically, as expected. Furthermore, the values of $\sigma_{\epsilon}$ and $\sigma_{\epsilon,2}$ are consistent with those obtained using the marginalized likelihood.
@@
Figure 14: Results of the joint FreeIC inference on Mock B using non-marginalized likelihood and Eulerian bias. We show the posterior for all noise parameters, bias parameters and $\beta$. Inference was performed at $\Lambda=0.15\,h\,{\rm Mpc}^{-1}$.
The same analysis using the non-marginalized likelihood has been applied to Mock A sampled with the Lagrangian bias. The results of this analysis are summarized in Fig. 15. Similarly as in the case of mock B, we find that the error bar $\sigma(\beta)$ is the same as when bias parameters are marginalized analytically.
Figure 15: Results of the joint FreeIC inference on Mock A using non-marginalized likelihood and Lagrangian bias. We show the posterior for all noise parameters, bias parameters and $\beta$. Inference was performed at $\Lambda=0.15\,h\,{\rm Mpc}^{-1}$.